# A Bayesian hierarchical mixture cure modelling framework to utilize multiple survival datasets for long-term survivorship estimates: A case study from previously untreated metastatic melanoma

### v1

@@ 3.1.1 Issues modelling multiple end-points
There are several potential issues with existing approaches to cure modelling, when applied to an HTA context in particular.
@@
Secondly, time to event data are typically right-censored due to administrative censoring or loss to follow up. The censoring times may be too early in order to adequately characterise a survival model for HTA for a single end-point. In particular, OS times are bounded below by PFS times by definition, so will be subject to more censoring at the same cut-point. By taking advantage of the set of end-points together, the data paucity may be alleviated to give better fits and extrapolation to the full lifetimes of patients. This is particularly important for health economics modelling in order to gauge the full cost and effect of an intervention.
Furthermore, if there is heavy censoring for all endpoints, such as early data cuts, and the number of endpoints is small, then the data alone may not represent the true underlying survival curves, including the plateau for the cure fractions. At this point, additional sources of information should be principally incorporated into the model. A Bayesian paradigm naturally allows for the inclusion of external evidence from sources such as historical data and expert opinion. It is reasonable to assume, in the absence of additional information, that the cure fractions for all end-points are drawn from the same or very similar distribution. In order to do this we can increase the amount of strength to borrow between end-point data which strikes a balance between the information (or lack of) in the data and other knowledge [16]. Practically, this corresponds to a reduction in the global cure fraction variance hyper parameter which in the limit is a base model with $\sigma_{k}=0$, equivalent to complete pooling of the data.
We will now present a Bayesian hierarchical mixture cure model for multiple end-points. This has the double advantage of borrowing information between end-points (e.g. in the case of OS and PFS, the likely more mature PFS data can inform the often highly censored OS analysis) and obtaining a single cure fraction estimate.
@@ 3.2 The hierarchical mixture cure model
Firstly, the approach taken so far, and presented in Section 3.1, is to model the cure fraction corresponding to each end-point completely separately (no pooling), assuming that they are independent.
@@
Secondly, we can assume that the cure fraction is the same for all end-points and so pool (or ”lump” together) the samples (complete pooling). In our example, PFS can be used as a proxy for OS since there are fewer missing data. There may be no reason why the cure fraction should be different between different end-points, especially if individuals are observed for long enough. In this case
@@
Lastly, and the approach proposed in this paper, is a compromise between the first two approaches called partial-pooling. We propose a hierarchical structure on the cure fraction assuming exchangeability between all end-points $j$ for each treatment $k$. We assume that there is a single, shared ”common” cure fraction from which the separate end-point specific cure fractions are obtained, corresponding to equation (2) but where $\pi_{k},\beta^{\pi}_{k}$ are no longer for particular end-points as before but now represent ”global” values. Using the global cure fraction we can define the second-level end-point specific cure fractions as
@@
where $\nu_{k}=\text{logit}(\pi_{k})$ and $\sigma_{k}^{2}$ is the between-group (end-points) variance. The end-point specific cure fractions in our example are denoted $\pi_{PFS},\pi_{OS}$. Figure 2 shows a directed acyclic graphical (DAG) representation of the Bayesian hierarchical mixture cure modelling framework for two end-points. The equivalent DAGs for the complete and no pooling models are given in the Appendix. Note that pooling refers to the cure fractions. The survival times are not pooled; the parameters for the uncured survival models ($\phi$, $\beta^{\lambda}$ in the latency model) are estimated using the separate sets of survival times for each end-point eg, in the complete pooling case we do not assume that PFS and OS have the same survival time distributions. However, in the complete and partial pooling case there is an indirect pooling effect for the latency models via the incidence model, so the fact that the cure fractions are similar/the same will influence the shapes of the survival curves.
@@
Figure 2: A hierarchical mixture cure model DAG for a trial with two end-points. The trial indices correspond to the motivating example end-points of PFS (1) and OS (2). Solid lines represent stochastic and dashed lines deterministic relationships, respectively. Cured patients have fixed hazards e.g. taken from life tables. The distribution of times for uncured patient regresses on covariates for the rate parameter and $\phi$ is the set of ancillary parameters.
4 Application
@@
This analysis was carried-out using the Stan inference engine [17] called from R v.4.3.1 [18] using the package rstan on a Windows 10 PC. Each fit employed 4 chains of 1000 iterations each, and a burn-in of 100 iterations. In order to determine convergence, we checked effective sample sizes, Markov chain standard error and performed posterior predictive checks. Details of the algorithm and model checking can be found in the Appendix. The code has been developed into an R package using a generalisable framework. This means that analyses for an arbitrary number of end-points can be easily performed with new data. The package is freely-available and can be downloaded from https://github.com/StatisticsHealthEconomics/multimcm.
@@ 4.1 Background survival
In many cure fraction models, it is common to have $S_{b}(t,x)=1$ for all $t$ under consideration, which is reasonable in the short-term. However, we are focused on the full life-course of an individual in an HTA context and so include a general population all-cause mortality competing risk. Thus, the uncured population is subject to a background mortality risk in addition to the risk from the event of interest.
We used the World Health Organisation (WHO) life tables by country for the latest year available of 2016 [19] to inform the background mortality in the mixture cure model. The baseline hazards are the expected mortality rate for each patient at the age at which they experience the event. The mortality data are adjusted for country, age and gender. Thus, this structure provides a granular account of the different patient profiles in the trial. The WHO reports conditional probabilities of death in 5-year intervals until age 85. A constant annual mortality rate is reported for individuals over 85. They assumed that no-one lives beyond 100 years. Further details are provided in the supplementary material.

### v2

@@ 3.1.1 Issues modelling multiple end-points
There are several potential issues with existing approaches to cure modelling, when applied to an HTA context, in particular.
@@
Secondly, time-to-event data are typically right-censored due to administrative censoring or loss to follow-up. The censoring times may be too early in order to adequately characterise a survival model for HTA for a single end-point. In particular, OS times are bounded below by PFS times by definition, so will be subject to more censoring at the same cut-point. By taking advantage of the set of end-points together, the data paucity may be alleviated to give better fits and extrapolation to the full lifetimes of patients. This is particularly important for health economic modelling to gauge the full cost and effect of an intervention.
Furthermore, if there is heavy censoring for all endpoints, such as early data cuts, and the number of endpoints is small, then the data alone may not represent the true underlying survival curves, including the plateau for the cure fractions. At this point, additional sources of information should be principally incorporated into the model. A Bayesian paradigm naturally allows for the inclusion of external evidence from sources such as historical data and expert opinion. It is reasonable to assume, in the absence of additional information, that the cure fractions for all end-points are drawn from the same or very similar distribution. In order to do this, we can increase the amount of strength to borrow between end-point data which strikes a balance between the information (or lack of) in the data and other knowledge 16. Practically, this corresponds to a reduction in the global cure fraction variance hyper-parameter which in the limit is a base model with $\sigma_{k}=0$, equivalent to complete pooling of the data.
We will now present a Bayesian hierarchical mixture cure model for multiple end-points. This has the double advantage of borrowing information between end-points (e.g., in the case of OS and PFS, the likely more mature PFS data can inform the often highly censored OS analysis) and obtaining a single cure fraction estimate.
@@ 3.2 The hierarchical mixture cure model
First, the approach taken so far, and presented in Section 3.1, is to model the cure fraction corresponding to each end-point completely separately (no pooling), assuming that they are independent.
@@
Secondly, we can assume that the cure fraction is the same for all end-points and so pool (or "lump" together) the samples (complete pooling). In our example, PFS can be used as a proxy for OS since there are fewer missing data. There may be no reason why the cure fraction should be different between different end-points, especially if individuals are observed for long enough. In this case
@@
Lastly, and the approach proposed in this paper, is a compromise between the first two approaches called partial-pooling. We propose a hierarchical structure on the cure fraction assuming exchangeability between all end-points $j$ for each treatment $k$. We assume that there is a single, shared "common" cure fraction from which the separate end-point specific cure fractions are obtained, corresponding to equation (2) but where $\pi_{k},\beta^{\pi}_{k}$ are no longer for particular end-points as before but now represent "global" values. Using the global cure fraction, we can define the second-level end-point specific cure fractions as
@@
where $\nu_{k}=\text{logit}(\pi_{k})$ and $\sigma_{k}^{2}$ is the between-group (end-points) variance. The end-point specific cure fractions in our example are denoted $\pi_{PFS},\pi_{OS}$. Figure 2 shows a directed acyclic graphical (DAG) representation of the Bayesian hierarchical mixture cure modelling framework for two end-points. The equivalent DAGs for the complete and no pooling models are given in the Appendix. Note that pooling refers to the cure fractions. The survival times are not pooled; the parameters for the uncured survival models ($\phi$, $\beta^{\lambda}$ in the latency model) are estimated using the separate sets of survival times for each end-point e.g., in the complete pooling case we do not assume that PFS and OS have the same survival time distributions. However, in the complete and partial pooling case, there is an indirect pooling effect for the latency models via the incidence model, so the fact that the cure fractions are similar/the same will influence the shapes of the survival curves. A more general DAG for any number of end-points would have an additional plate (box) around $\pi_{jk}$ with a single individual level model nested within it.
@@
Figure 2: A hierarchical mixture cure model DAG for a trial with two end-points. The trial indices correspond to the motivating example end-points of PFS (1) and OS (2). Solid lines represent stochastic and dashed lines deterministic relationships, respectively. Cured patients have fixed hazards, e.g. taken from life tables. The distribution of times for uncured patient regresses on covariates for the rate parameter and $\phi$ is the set of ancillary parameters.
4 Simulation study
To assess the performance of our method in addressing the issues set out above, we conducted a simulation study and compared the outcome between defined scenarios 17. We followed the structured approach for designing and reporting simulation studies, described in 18, which involves defining aims, data-generating mechanisms, estimands, methods, and performance measures (“ADEMP”).
4.1 Scenarios and data generating process
We simulated survival times from various hierarchical mixture cure models. We decided to do this rather than resampling from an existing data set because of the flexibility and extended scenarios this made available. For example, the real data example in Section 5 would be constrained to only two end-points.
The target estimates used were the restricted mean survival time (RMST) and cure fractions ($\pi$). The median survival time ($t_{0.5}$) was not used. These were chosen because they are particularly relevant for health economic evaluation. An alternative approach would have been to use, e.g. the Concordance Index (C-index) or log-rank test more favoured in the field of survival analysis. We shall use a particular version of RMST appropriate for a mixture cure model [^1] In a similar way, the median survival time would be adjusted for the cure fraction model. Where the median time for the uncured fraction is that $t$ such that $S_{u}(t)=0.5$, for the entire population, we require $S(t)=0.5$. If we assume that there is no additional background mortality, then $t_{0.5}$ is the solution of $S_{u}(t)=(0.5-\pi)/(1-\pi)$.. Defining $\tau$ as the maximum (cut-off) time then
$RMST=\pi\tau+(1-\pi)\sum_{0}^{\tau}S(t)\;\text{d}t.$
The underlying data generating model is of the form
$\text{logit}(\pi_{j})$ $\sim\text{N}(\mu_{\text{true}},\sigma^{2}_{\text{true}})$
$s_{ij}$ $\sim\text{Bern}(\pi_{j})$
$t^{u}_{ij}$ $\sim\text{Weibull}(k_{j},\lambda_{j})$ (4)
$t_{ij}^{\text{obs}}$ =$=\begin{cases}t^{u}_{ij},&\text{if }s_{ij}=0\text{ (not cured)}\\ \tau,&\text{if }s_{ij}=1\text{ (cured)},\end{cases}$
where event time $t_{ij}^{\text{obs}}$ for individual $i$ and end-point $j$ are generated from a Weibull distribution with the only censoring at time $\tau$ due to being cured. No covariates were included. An individual’s cured status $s_{ij}$ is generated from a Bernoulli distribution with probability $\pi_{j}$ sampled from a global cure fraction Normal distribution on the logit linear scale with parameters $\mu_{\text{true}},\sigma^{2}_{\text{true}}$. We don’t include any censoring in the data generation because this has a similar effect to sample size for our purposes and would further increase the number of simulation scenarios. That is, introducing censoring is a form of information loss which reduces the effective sample size (although still providing partial information).
The performance measures calculated were bias, empirical standard error (empSE) and coverage. Bias is the difference between the expected value of an estimator and the true value. The empSE is the true variability of the estimator. Finally, the coverage probability is the probability that the credible intervals contain the true value which we use the 95% level. We also included relative bias (RB), which is the bias scaled by the true value, because different values of $\pi_{j}$ are used for different data-generating mechanisms and this permits a more straightforward comparison. Target mean estimates $\widehat{RMST}$ and $\widehat{\pi}$ were obtained from posterior samples.
The underlying true values that were held fixed between scenarios were selected as follows. The between-group cure fraction standard deviation was held fixed because the simulation study is concerned with the relation with the cure fraction priors. We also make the assumption that it is narrow enough to justify a hierarchical model and exchangeability. Note again that $n_{sim}$ can be thought of in a sense as an equivalent larger sample size in the presence of censoring. The true global cure fraction was 0.2; the global cure fraction standard deviation was assigned a zero-truncated Cauchy of $\text{Cauchy}^{+}(0.4,2.4)$ (corresponding to mean 19, median 2.4 and 95% interval 0.1 - 55); for the separate cure fraction model, the cure fraction on the logistic scale was assigned the prior distribution $\text{N}(-1,(0.2+0.4)^{2})$ to correspond to the hierarchical specification; cut-point time $\tau={5}$; the true latent survival model distributions were a 1:1 ratio of $\text{Weibull}(k=1,\lambda=1)$ and $\text{Weibull}(k=1,\lambda=4)$. This was chosen to ensure that half of the curves had not converged sufficiently to the plateau at the cure fraction, and thus allowed for the potential borrowing of information.
The factors to vary were chosen over two levels of model inputs, corresponding to high or low values. These factors were varied fully factorially. For the purposes of clarity and space, a selection of these will be presented here, and the full set of plots and tables is presented in the Appendix. We shall reference items in the Appendix with a preceding A. In total, the analysis consists of 32 scenarios as follows.
In terms of the data simulation, the endpoints, i.e. number of survival curves, was $n_{e}={3,10}$, and sample size was $n_{sim}={10,100}$. In terms of the model prior distributions, the latent survival curves were given informative priors ${k\sim\text{Gamma}(1000,1000)}$, and either $\lambda\sim\text{logN}(1.4,0.002)$ or $\lambda\sim\text{logN}(0,0.01)$; and weak priors $k\sim\text{Gamma}(1000,100)$, and either $\lambda\sim\text{logN}(1.4,0.02)$ or $\lambda\sim\text{logN}(0,0.1)$. The former corresponds to the situation when we have good information about the underlying (uncured) curves but perhaps not for the cure fractions. Similarly, the global cure fraction scenarios are for when the global mean is strongly centred at the true value or weak centred at 0.5 [95% interval: 0.2 - 0.8], indicating little prior knowledge; the mean global cure fraction priors on the logit scale were either informative $\beta^{\pi}\sim N(-1,0.01)$ or weak $\beta^{\pi}\sim N(0,0.7)$. Finally, the between-group cure fraction prior standard deviation was assigned an informative prior $\sigma\sim\text{logN}(0.05,0.05)$ and a weak prior $\sigma\sim\text{logN}(0.05,0.1)$. The full list of scenarios is presented in Table A18 in the Appendix. Finally, note that the varying of the different prior distributions is a type of robustness analysis, from which we can investigate the influence of the different choices. Having the global cure fraction prior centred at 0.5 in our analysis is an example of what’s called an ‘optimistic’ prior, i.e. bias upward.
4.2 Simulation results
The main analysis produced 1000 posterior samples for each scenario data set, which were replicated 1000 times following equation (4). The runtime exceeded several hours using a Myriad high-performance cluster. Additional figures and tables of results found in the Appendix include histograms of posterior distributions of the performance measures (Figures A18 and A19), lollipop plots comparing separate and hierarchical results (Figures A20-A25), and zip plots of coverage (Figures A26-A34). For an overall summary, Table 1 shows the performance measure estimates averaged over all curves within scenarios. Figures A22-23 in the Appendix give corresponding heat maps for easy pattern recognition.
RMST Cure Fraction
Bias Coverage empSE RB Bias Coverage empSE RB
Scenario Hier Sep Hier Sep Hier Sep Hier Sep Hier Sep Hier Sep Hier Sep Hier Sep
1 0.01 0.01 0.99 0.86 0.03 0.00 0.00 0.00 0.00 0.00 0.99 0.84 0.01 0.00 0.01 0.00
2 0.01 0.01 0.98 0.85 0.03 0.00 0.00 0.00 0.00 0.00 0.98 0.83 0.01 0.00 0.01 0.00
3 0.01 0.01 0.96 0.84 0.03 0.01 0.00 0.00 0.00 0.00 0.96 0.84 0.01 0.00 0.00 0.00
4 0.01 0.01 0.94 0.85 0.03 0.01 0.00 0.00 0.00 0.00 0.93 0.84 0.01 0.00 0.00 0.00
5 0.57 0.50 0.21 0.10 0.22 0.16 0.18 0.16 0.04 0.00 0.95 0.83 0.08 0.00 0.15 0.01
6 0.55 0.42 0.34 0.14 0.28 0.21 0.19 0.14 0.06 0.00 0.94 0.83 0.09 0.00 0.22 0.01
7 0.57 0.39 0.12 0.02 0.14 0.09 0.19 0.14 0.12 0.01 0.35 0.72 0.05 0.01 0.46 0.05
8 0.52 0.37 0.18 0.04 0.15 0.10 0.18 0.14 0.10 0.01 0.48 0.75 0.05 0.00 0.36 0.04
9 0.01 0.01 1.00 0.97 0.06 0.01 0.00 0.00 0.00 0.00 1.00 0.97 0.02 0.00 0.01 0.00
10 0.01 0.01 0.99 0.97 0.06 0.01 0.01 0.00 0.00 0.00 0.99 0.97 0.02 0.00 0.01 0.01
11 0.01 0.01 0.98 0.97 0.04 0.02 0.00 0.00 0.00 0.00 0.98 0.97 0.02 0.01 0.00 0.00
12 0.01 0.01 0.97 0.97 0.04 0.03 0.00 0.00 0.00 0.00 0.97 0.97 0.01 0.01 0.00 0.00
13 0.61 0.51 0.22 0.13 0.23 0.16 0.20 0.16 0.06 0.00 0.96 0.97 0.09 0.00 0.21 0.02
14 0.57 0.43 0.37 0.20 0.27 0.22 0.20 0.14 0.07 0.00 0.94 0.97 0.09 0.00 0.27 0.02
15 0.57 0.42 0.12 0.04 0.14 0.09 0.19 0.14 0.12 0.03 0.35 0.65 0.05 0.01 0.47 0.11
16 0.51 0.39 0.18 0.06 0.15 0.10 0.18 0.15 0.10 0.02 0.48 0.73 0.05 0.01 0.36 0.09
17 0.20 0.36 0.93 0.92 0.19 0.20 0.08 0.13 0.07 0.13 0.94 0.93 0.07 0.07 0.26 0.50
18 0.08 0.38 0.95 0.91 0.15 0.23 0.03 0.16 0.02 0.13 0.95 0.92 0.05 0.07 0.09 0.48
19 0.03 0.08 0.95 0.94 0.10 0.14 0.01 0.03 0.01 0.03 0.95 0.94 0.04 0.05 0.03 0.10
20 0.01 0.08 0.94 0.94 0.06 0.15 0.01 0.03 0.00 0.03 0.94 0.95 0.02 0.05 0.01 0.10
21 0.91 0.90 0.09 0.18 0.19 0.21 0.32 0.31 0.20 0.21 0.34 0.54 0.08 0.08 0.75 0.78
22 0.84 0.87 0.10 0.26 0.21 0.24 0.32 0.33 0.17 0.20 0.17 0.60 0.05 0.09 0.65 0.74
23 0.63 0.62 0.06 0.08 0.15 0.14 0.22 0.21 0.14 0.15 0.23 0.30 0.05 0.05 0.53 0.56
24 0.56 0.57 0.11 0.13 0.15 0.15 0.21 0.21 0.11 0.12 0.44 0.45 0.05 0.05 0.41 0.45
25 0.19 0.36 0.95 0.91 0.20 0.21 0.07 0.13 0.07 0.13 0.95 0.92 0.07 0.07 0.25 0.50
26 0.08 0.38 0.97 0.91 0.15 0.23 0.03 0.15 0.02 0.13 0.97 0.92 0.05 0.07 0.09 0.48
27 0.03 0.08 0.97 0.95 0.10 0.14 0.01 0.03 0.01 0.03 0.97 0.95 0.04 0.05 0.03 0.10
28 0.01 0.08 0.96 0.94 0.06 0.15 0.00 0.03 0.00 0.03 0.96 0.95 0.02 0.05 0.01 0.10
29 0.91 0.91 0.11 0.16 0.20 0.20 0.32 0.32 0.20 0.21 0.39 0.54 0.08 0.09 0.74 0.78
30 0.84 0.87 0.13 0.26 0.22 0.25 0.32 0.33 0.17 0.20 0.25 0.60 0.06 0.09 0.65 0.73
31 0.63 0.62 0.07 0.09 0.14 0.14 0.22 0.21 0.14 0.15 0.25 0.31 0.05 0.05 0.54 0.56
32 0.56 0.57 0.11 0.13 0.15 0.15 0.20 0.21 0.11 0.12 0.44 0.45 0.05 0.05 0.41 0.45
Table 1: Performance measures estimates for the hierarchical (Hier) and separate (Sep) models giving restricted mean survival time (RMST) and cure fraction ($\pi$) results. Values are averaged over all curves within scenarios.
4.2.1 Performance of the target estimates
For both the RMST and cure fraction the picture is similar. If the uncured survival curve and cure fraction prior distributions are informative then the separate and hierarchical models perform well with comparably small measures for bias, irrespective of whether an informative or weak prior is placed on the between group variance. However, in these scenarios the coverage is better for the separate model. Alternatively, when both uncured survival and cure fraction have only weakly informative priors then both separate and hierarchical models demonstrate poor bias, irrespective of the between group prior (scenarios 1-4, 9-12). The scenarios of note are when we have an informative prior for the uncured survival curve parameters but only weak prior knowledge for the cure fraction (scenarios 17-20, 25-28). In this case, the hierarchical model bias is clearly better than for the separate model. The coverage is consistently better for scenarios with an informative survival curve prior too. In term of sample size, the biggest difference between small and large samples in for when the cure fraction has a weak prior; or alternatively, with strong prior evidence then the measures are less sensitive to sample size. For coverage, when the bias is relatively large the coverage is low and vice-versa. In particular, for a small number of end-points and small sample size, the hierarchical model performs better in terms of bias when an informative survival curve prior is used with a weak cure fraction prior (scenarios 18 and 26). Conversely, for a weak prior for uncured survival and informative cure fraction (scenarios 5-8, 13-16) the coverage and bias are worse for the hierarchical model than the separate model, but in general when the uncured survival prior is weak the coverage is poor.
In summary, the pattern of the performance measures between target estimates is similar. This is likely because the RMST is dependent on the cure fraction of the underlying curve. Overall, the simulation study shows that performance of estimates can vary depending on the cure fraction, uncured survival curves and the sample size. The scenarios when there is less prior knowledge about the uncured survival curve and more information about the cure fraction seem to preference the separate model, whereas conversely scenarios where the hierarchical model is most beneficial include those with informative priors on the uncured survival and uncertainty about the cure fraction. This latter case seems reasonable in reality when, for instance, a new drug impacts an uncertain proportion of long-term survivors for a well understood underlying process. In contrast, in other situations, it may not be possible to obtain direct evidence about an untreated population due to ethical considerations.
5 Real data example
@@
This analysis was carried-out using the Stan inference engine 19 called from R v.4.3.1 20 using the package rstan on a Windows 11 PC with 12 cores. Each fit employed the default of 4 chains of 2000 iterations each, and a warm up of 100 iterations. In order to determine convergence, we checked effective sample sizes, Markov chain standard error and performed posterior predictive checks. Details of the algorithm and model checking can be found in the Appendix. The code has been developed into an R package using a generalisable framework. This means that analyses for an arbitrary number of end-points can be easily performed with new data. The package is freely-available and can be downloaded from https://github.com/StatisticsHealthEconomics/multimcm.
@@ 5.1 Background survival
In many cure fraction models, it is common to have $S_{b}(t,x)=1$ for all $t$ under consideration, which is reasonable in the short-term. However, we are focused on the full life-course of an individual in an HTA context and so include a general population all-cause mortality competing risk. Thus, the uncured population is subject to a background mortality risk in addition to the (excess) risk from the event of interest.
We used the World Health Organisation (WHO) life tables by country for the latest year available of 2016 21 to inform the background mortality in the mixture cure model. The baseline hazards are the expected mortality rate for each patient at the age at which they experience the event. The mortality data are adjusted for country, age and gender. Thus, this structure provides a granular account of the different patient profiles in the trial. The WHO reports conditional probabilities of death in 5-year intervals until age 85. A constant annual mortality rate is reported for individuals over 85. The life table is produced assuming that no-one lives beyond 100 years. Further details are provided in the Appendix.